# Lagrangian Motion Fields for Long-term Motion Generation

### v1

@@ 4.3 Qualitative Evaluation and User Study
TABLE IV: Ablation Studies on FineDance dataset. “#Cls” denotes the number of clusters.
@@
EDGE (6D) - - 0.211 20.04% 94.34 50.38
Ours (3D) 1000 w/o 0.201 21.17% 57.81 36.37
Ours (3D) 1000 w 0.219 23.16% 53.79 33.84
Ours (3D) 2000 w 0.235 24.91% 50.27 31.63
Ours (6D) 2000 w 0.228 12.63% 55.49 24.93
TABLE V: Ablation Studies on HumanML3D dataset.
@@
4.4 Ablation Study
We conduct ablation studies on different coordinate systems, cluster numbers and the proposed coherent loss to evaluate the robustness and generalizability of our method.
@@ 5 Applications
Duration-Controlled Motion Generation. In many scenarios, such as virtual storytelling, cinematic animation, or fitness applications, controlling the duration of a motion sequence is essential. By leveraging our supermotion representation, which inherently encodes duration information, duration control can be achieved by applying the standard masked denoising technique described by Tseng et. al. [9]. Given a duration condition $D$, we decompose it into a sequence $[d_{0},d_{1},\dots,d_{M-1}]$ such that $d_{min}\leq d_{i}\leq d_{max}$ and $\sum_{s=0}^{M-1}d_{s}=D$. During denoising, we update the duration component of the supermotion as follows:
@@ 6 Discussion and Conclusion
While our method offers significant advantages, a common challenge with abstract representations like supermotions, similar to superpixels and supervoxels, is the potential for oversmoothing in motion sequences, which can slightly affect realism. To mitigate this, we have implemented a lightweight refinement module that effectively enhances motion details. In future work, we plan to integrate more advanced refinement techniques and explore adaptive mechanisms to better capture intricate motion dynamics. Looking ahead, we also aim to extend our approach to additional applications, such as real-time motion generation and interactive virtual environments, to further validate its versatility and robustness.

### v2

@@ 4.4 User Study
4.5 Ablation Study
In this section, we evaluate the key components of our framework, including different coordinate systems, cluster numbers, the proposed coherence loss, and the refinement module, to assess its robustness and generalizability.
TABLE IV: Ablation Studies on FineDance dataset. “#Cls” denotes the number of clusters. “Refine.” indicates whether the Refinement Module is used. “Params” shows the number of parameters in millions (M) for the Refinement Module.
@@ 4.5 Ablation Study
EDGE (6D) - - - - 0.211 20.04% 94.34 50.38 8.13 6.45
Ours (3D) 1000 w/o 60.78M 0.201 21.17% 57.81 36.37 5.21 4.98
Ours (3D) 1000 w 60.78M 0.219 23.16% 53.79 33.84 5.35 5.08
Ours (3D) 2000 w 60.78M 0.235 24.91% 50.27 31.63 5.63 5.27
Ours (6D) 2000 w 60.78M 0.228 12.63% 55.49 24.93 5.46 5.82
Ours (6D) 2000 w 41.86M 0.225 12.89% 56.82 25.74 5.83 5.76
Ours (6D) 2000 w - 0.202 36.15% 88.60 40.52 8.17 5.80
TABLE V: Ablation Studies on the HumanML3D dataset.
@@
Refinement Module. To assess the refinement module’s impact, we compare performance with and without it in our ablation study. As shown in Table IV, removing refinement (✗under “Refine.”) degrades motion quality. Although the ablation variant achieves higher numerical diversity scores, these gains primarily result from artifacts such as excessive foot sliding and incoherent transitions, which introduce noisy variance rather than meaningful motion differences. In contrast, incorporating the refinement module improves temporal coherence and realism, thereby suppressing spurious variance and substantially enhancing overall quality while preserving perceptual diversity.
Moreover, the module is parameter-efficient, as performance remains stable even with fewer parameters. This underscores the supermotion representation’s role in providing structural guidance, enabling efficient quality improvement with minimal overhead.
@@ 5 Applications
Duration-Controlled Motion Generation. In many scenarios, such as virtual storytelling, cinematic animation, or fitness applications, controlling the duration of a motion sequence is essential. By leveraging our supermotion representation, which inherently encodes duration information, duration control can be achieved by applying the standard masked denoising technique described by Tseng et. al. [9]. Given a duration condition $D$, we decompose it into a sequence $d_{0},d_{1},\dots,d_{M-1}]$ such that $d_{min}\leq d_{i}\leq d_{max}$ and $\sum_{s=0}^{M-1}d_{s}=D$. During denoising, we update the duration component of the supermotion as follows:
@@ 6 Discussion and Conclusion
Figure 12: Limitations of our method: (a) Rapid directional changes, leading to frame-to-frame discontinuities. (b) High-frequency repetitive actions, where distinct patterns are merged.
While our method offers significant advantages, abstract representations like supermotions, similar to superpixels and supervoxels, can lead to oversmoothing in motion sequences, slightly affecting realism. Although the proposed lightweight refinement module enhances motion details, it struggles with rapid directional changes and high-frequency localized actions, where abrupt transitions and fine-grained motion nuances are not fully preserved, as shown in Fig. 12. To address these limitations, we plan to explore more advanced refinement techniques and adaptive mechanisms to better capture intricate motion dynamics. Furthermore, we aim to extend our approach to real-time motion generation and interactive virtual environments, further validating its versatility and robustness.